OpenStudy (anonymous):

two circles have radii of 10 and 13 what is their ratio of their circumferences?

OpenStudy (anonymous):

C = 2pi r 20pi to 26pi 26/20 = 1.3 to 1

OpenStudy (anonymous):

the options are a. 100/169 b. 10/13 c. sqrt130/13

OpenStudy (anonymous):

@MattBrotz

OpenStudy (anonymous):

b

OpenStudy (anonymous):

same thing as 1.3/1

OpenStudy (anonymous):

@MattBrotz can you help me with some other stuff?

OpenStudy (anonymous):

\[Circumference = 2*\pi*r\] \[Circumference_{r=10} : Circumference_{r=13}\] \[2*\pi*10 : 2*\pi*13\] \[10 : 13\]

OpenStudy (anonymous):

@madddiepaige You may want to fill out a new question for other questions.